OpenStudy (anonymous):

I gotta go to a wedding in like less than 16 mins and I'm stuck on homework. :( HELP PROVE: (cscx(cosx+sinx)^2)/1+sin2x = cotx/cscx

OpenStudy (anonymous):

\[\frac{ \csc x \left( \cos x+\sin x \right)^2 }{1+\sin 2x }=\frac{ \csc x \left( \cos ^2x+\sin ^2x+2\sin x \cos x \right) }{ 1+\sin 2x }\] \[=\frac{ \csc x \left( 1+\sin 2x \right) }{ 1+\sin 2x }=\csc x\] \[\frac{ \cot x }{ \csc x }=\frac{\cos x }{ \sin x }*\frac{ \sin x }{ 1 }=\cos x\] check your statement.

OpenStudy (anonymous):

I was give this question but i would always get cscx.

OpenStudy (anonymous):

left hand side =csc x right hand side=cosx

OpenStudy (anonymous):

I guess it can't be proven.

OpenStudy (anonymous):

last question but, do you know how to solve for cotx - (cos2x/sinxcosx) = tanx

OpenStudy (anonymous):

can you rewrite the equation to prove it?

OpenStudy (zehanz):

Seems you're late for the wedding now. Hurry!